OpenStudy (anonymous):

sec(tan^-1 u + cos^-1 v) sec(tan^-1 u + cos^-1 v) @Mathematics

OpenStudy (anonymous):

ick. use addition angle formula for cosine and then take the reciprocal

OpenStudy (jamesj):

oh, I just posted on this the second time you asked (@physics)

OpenStudy (jamesj):

Here's a copy ========= well, cos(x + y) = cos x . cos y - sin x . sin y Hence cos(tan^-1 u + cos^-1 v) = cos(arctan u).v - sin(arctan u).sin(arccos v) You should be able to reduce that expression further now (if you're stuck, draw a triangle). Then take the reciprocal to get back to sec.

OpenStudy (anonymous):

\[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\] with \[a=\tan^{-1}(u)\] \[b=\cos^{-1}(v)\] so you need \[\cos(\tan^{-1}(u))=\frac{1}{\sqrt{u^2+1}}\] \[\cos(\cos^{-1}(v))=v\] \[ sin(\tan^{-1}(u))=\frac{u}{\sqrt{u^2+1}}\] \[\sin(\cos^{-1}(v))=\sqrt{1-v^2}\] and plug all that mess in

OpenStudy (anonymous):

what jamesj said. took me like ten minutes to type this!